OpenStudy (anonymous):

4 math questions?medal&fan! Which point is on the graph of y=3/x? a. (6, 2) b. (6, 1/2 ) c. (0, 3) d. (3, 0)? "Four less than three times a number" can be written as 4 - 3n . a. true b. false Which answer best describes the relationship between the two lines defined by y = -2x + 9 and y = ½x - 3 ? a. Parallel b. Perpendicular c. The lines intersect but they are neither parallel or perpendicular. d. The two lines never intersect and are NOT parallel. Solve the linear system. You may use any method. 3x + 2y = 14 and 5x - 2y = 2 Your answer should be an ordered pair. EX(5,6)

OpenStudy (anonymous):

1.d 2.true 3.d 4. the first one is (14/3,7) nd second one is (2/5,-1)

OpenStudy (anonymous):

Which point is on the graph of y = 3/x? B: (6,1/2) "Four less than three times a number" can be written as 4 - 3n B: False Which answer best describes the relationship between the two lines defined by y = -2x + 9 and y = ½x - 3 ? C: The lines intersect but they are neither parallel or perpendicular Solve the linear system. You may use any method. 3x + 2y = 14 and 5x - 2y = 2 Your answer should be an ordered pair. EX(5,6) 3x + 2y = 14 2y = 14 - 3x y = 7 - 3/2x 5x - 2y = 2 5x - 2(7-3/2x) = 2 5x -14 + 3x = 2 8x - 14 = 2 8x = 16 x = 2 3(2) + 2y = 14 6 + 2y = 14 2y = 8 y = 4 Answer: (2,4)